OpenStudy (anonymous):

What is the x-value of the solution to y=-2x+10 and y=2/3x+2. A:1 B:2 C:3 D:4

OpenStudy (anonymous):

both equations are expressing y , so they must be equal since y=y, so equate them and simplify

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

y = -2x +10 y = 2/3x + 2 don't they both express y

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so this means that: 2/3x +2 = -2x +10,

OpenStudy (anonymous):

ok so whats the answer that you would choose?

OpenStudy (anonymous):

which answer did you get?

OpenStudy (anonymous):

oh hold on ima try it out

OpenStudy (anonymous):

idk but its giving me decimals

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

bring like terms together like this: 2/3x +2x = 10-2............remember when the -2x crosses the = sign it becomes positive. 2 2/3x = 8 .......... change into improper fraction 8/3x = 8........... multiply both sides by reciprocal i.e 3/8 x = 8 * 3/8 = 3, any where you don't follow?

OpenStudy (anonymous):

its three i got it!!!!!! :) dude you rock

OpenStudy (anonymous):

you right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you welcome